1,561 views
0 0 votes
Consider a hash table with 10 slots.

Collisions are resolved using linear probing. The probability that if first 3 slots are unfilled and 4th insertions leads to a collisions ________.

(Upto 3 decimal places)

 

// my answer is 7/10 * 6/10 * 5/10 * 3/7 = 90/1000 = 0.09

Please log in or register to answer this question.

Position:
Show:

No related questions found